What is the molarity of h2so4 solution that has a density 1.84 g/cc at 35` Celsius and contain 98% h2so4 by weight

Grade:11

1 Answers

Vikas TU
7 years ago
Volume = mass/densiyt = > 98/1.84 = > 53.26 ml. of soln.
Moles of H2SO4 = > 98/98 = 1 mol.
Molarity = moles/volume in ltres. = > 1/(56.26/1000) = > 1000/56.26 = > 17.77 M.